2x-y =6 x + y = -3
Answer by mananth(16946)

2 x -1 y = 6 .............1
1 x + 1 y = -3 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
2 x -1 y = 6
1 x + 1 y = -3
Add the two equations
3 x = 3
/ 3
x = 1
plug value of x in (1)
2 x -1 y = 6
2 -1 y = 6
-1 y = 6 -2
-1 y = 4
y = -4
(1,-4)